OpenStudy (anonymous):

What is the perimeter of a regular octagon with a side length of 8 cm?

OpenStudy (anonymous):

8 * 6 = 48 cm

OpenStudy (anonymous):

side=8cm perimeter=8a=8*8=64

OpenStudy (anonymous):

sorry - hakkun is right an octagon has 8 sides!!